OpenStudy (anonymous):

The volume of a cone is 300 cubic kilometers. If the radius of the base is 5 kilometers, what is the height of the cone? Use 3.14 for pi Show your work and round to the nearest tenth.

OpenStudy (anonymous):

V = 1/3 pi r^2 h substitute 300 = 1/3 (3.14)(6)^2 h solve for h 300 = 37.68h h = 8.0 km So the hight will be 8.0km

OpenStudy (anonymous):

V = 1/3 pi r^2 h substitute 300 = 1/3 (3.14)(6)^2 h solve for h 300 = 37.68h h = 8.0 km So the hight will be 8.0km

OpenStudy (anonymous):

V = 1/3 pi r^2 h substitute 300 = 1/3 (3.14)(6)^2 h solve for h 300 = 37.68h h = 8.0 km So the hight will be 8.0km